OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time (in years). Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Pleas

OpenStudy (anonymous):

so for both payments you will have to try and calculate it for the same period lets say one month. Do you know how to solve it now?

OpenStudy (anonymous):

No not really :(

OpenStudy (anonymous):

Payment Option 1: $16,000 - $3,000 = $13,000 after down payment $13,000 / 60months = $216/month Payment Option 2: $16,000 * 1.02 / 72months = $226.67/month

OpenStudy (anonymous):

i understand it now ....thanks

OpenStudy (anonymous):

ok, let me explain. Payment option 1: So we pay $3000 as a deposit first hand and then we pay the remainder in monthly payments so it will be calculated as below. 160000-3000=$13000 (total price we need to pay after deposit) 13000/60 = $216.67 (monthly payment)

OpenStudy (anonymous):

wait, u got the right idea and working out by shouldn't ur rate be 0.01?

OpenStudy (anonymous):

what would the rate be? its 1%

OpenStudy (anonymous):

1%=0.01

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

$16,000 * 0.01 / 72months = $235.56/month so payment to will still cost more.

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

on that site the P given is 9.5% and in the working out they show 0.095 as the rate

OpenStudy (anonymous):

no problem, medal thanks :P

OpenStudy (anonymous):

okay thanks for the link also